Home‎ > ‎Crew 162‎ > ‎

0126-GreenHabReport

Crew 162 GreenHab Officer Report Sol 03
Renee Garifi
1/26/16

Watered the bean and radish sprouts that were repotted yesterday with
~200 mL water. Took photos and will forward to Crew 161 who started the
plants.

Started watering process for the two experimental plant growth hardware
sets in the temporary grow tent. The active watering system had a data
logger leg that shook loose in shipping to Mars. Photos and description
sent to PI for feedback. Will be troubleshooting and watering the
cabbage seedlings tonight. Photos will be transmitted back to Earth for
analysis by the PI.

No issues encountered with any of the three plant experiments.